1. What is the magnetic field at y = -3.6cm. ______mT k?

6. What is the magnetic field at y = 0. ______mT k?

9. What is the magnetic field at y =+3.8. ______mT k?

12. What is the magnetic field at y = +9.5 cm. _____mT k?

The current in the wire at y =-6.0 cm is in the-X direction and the current in the wire at y = +6.0 cm is in the +x direction. The current in each wire is 25 A. Find the magnetic field at the following points on the y axis. y--6.0 cnm y=+6.0 cm y
Answer #1

Magnetic field at a distance 'r' from the current carrying wire is given by, B = μoI/2πr

At y = -3.6 cm, B = -(4π * 10-7 * 25 / 2π)[1/(0.06 - 0.036) + 1/(0.06 + 0.036)] k = -0.26 mT k

At y = 0, B = -(4π * 10-7 * 25 / 2π)[1/0.06 + 1/0.06] k= -0.17 mT k

At y = 3.8 cm, B = -(4π * 10-7 * 25 / 2π)[1/(0.06 + 0.038) + 1/(0.06 - 0.038)] k = -0.28 mT k

At y = 9.5 cm, B = (4π * 10-7 * 25 / 2π)[-1/(0.06 + 0.095) + 1/(0.095 - 0.06)] k = 0.11 mT k